OpenStudy (anonymous):

what is the center of the circle with this equation x2+y2+14x-12y=-69

OpenStudy (rock_mit182):

i guess you have to complete squares

OpenStudy (rock_mit182):

x^2 + 14x + 7^2 + y^2 -12y + 6 ^ 2 = -69 + 7^2 + 6^2 since we've already completed squares, we could factor this out

OpenStudy (rock_mit182):

(x+7)^2 + (y-6)^2 = -69 +49 + 36

OpenStudy (rock_mit182):

@lovelaced do you get it ? now you can solve this